Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability Help!!!

Expert replies
by anuu » Thu Apr 07, 2011 11:23 am
A circular target of unit radius is divided into four annular zones with outer radii 1/4, 1/2, 3/4, and 1, respectively. Suppose ten shots are fired independently and at random into the target.

a) What is the prob. that at most three shots land in the zone bounded by the circles of radius 1/2 and 1.
b) If five shots land inside the disk of radius 1/2, find the probability that at least one is in the disk of radius 1/4.

I'm trying to understand the logic behind this problem. I haven't arrived at a right concept yet.

Any suggestions pls???

Anuu
Join the discussion
Source: — Problem Solving |

by manpsingh87 » Fri Apr 08, 2011 12:52 am
anuu wrote:A circular target of unit radius is divided into four annular zones with outer radii 1/4, 1/2, 3/4, and 1, respectively. Suppose ten shots are fired independently and at random into the target.

a) What is the prob. that at most three shots land in the zone bounded by the circles of radius 1/2 and 1.
b) If five shots land inside the disk of radius 1/2, find the probability that at least one is in the disk of radius 1/4.

I'm trying to understand the logic behind this problem. I haven't arrived at a right concept yet.

Any suggestions pls???

Anuu
its a question of geometrical probability.!!!
1)
the region bounded by the circle of radius of 1 and the circle of 1/2=pi(1^2)-pi(1/2)^2 =pi3/4;
therefore probability of hitting in the bounded region=pi(3/4)/pi1= 3/4;

similarly probability of not hitting in the bounded region = 1-3/4=1/4;

as maximum of 3 shots can hit the bounded region, therefore here 4 cases are possible.
1) when no shot hits the target.
2) when 1 shot hits the target.
3) when 2 shots hit the target.
4) when 3 shots hit the target.

(1/4)^10+10C1*(3/4)*(1/4)^9+10C2*(3/4)^2*(1/4)^8+10C3(3/4)^3*(1/4)^7..!!

please try 2) i'll post my approach for 2 later..!!!!
O Excellence... my search for you is on... you can be far.. but not beyond my reach!
Join the discussion

by Ian Stewart » Fri Apr 08, 2011 5:14 am
anuu wrote:A circular target of unit radius is divided into four annular zones with outer radii 1/4, 1/2, 3/4, and 1, respectively. Suppose ten shots are fired independently and at random into the target.

a) What is the prob. that at most three shots land in the zone bounded by the circles of radius 1/2 and 1.
b) If five shots land inside the disk of radius 1/2, find the probability that at least one is in the disk of radius 1/4.
manpsingh's explanation above is excellent. These questions are nothing at all like real GMAT questions, however. For one thing, you'd never need to know what 'annular zones' are on the GMAT, and for another, the computation involved is too cumbersome for the test. I imagine many people will not understand even what the question is asking because of the language used. The question is describing something a bit like a dartboard: we have four concentric circles with radii 1/4, 1/2, 3/4 and 1.

What you may, on rare questions, need to understand is how to calculate geometric probabilities. If we take a much simpler setup:

A circle C of radius 1 is contained within a circle D of radius 2. If a random point is selected from within circle D, what is the probability this point is contained in circle C?

this is just a complicated way of asking for the ratio of the area of C to the area of D (the answer is (Pi*(1)^2)/(Pi*(2)^2) = 1/4). That is, geometric probability questions are really just asking for ratios of areas.

While it's not something you'll see on the GMAT, we can apply this to the second question:

If five shots land inside the disk of radius 1/2, find the probability that at least one is in the disk of radius 1/4.

Here we are asked the probability that 'at least one' is in the disk of radius 1/4. There would be several cases to consider if we did the problem directly (exactly one, exactly two, exactly three, exactly four and exactly five shots in the smaller circle). Because that looks like a lot of work, it's a good idea to reverse the problem: find the probability that NO shots are in the smaller circle, and then subtract this from 1. So the probability that at least 1 shot is in the smaller circle is equal to:

1 - Probability(0 shots are in the small circle)

Now, the probability that one shot is in the small circle is just equal to the ratio of the area of the small circle to the larger circle:

Pi*(1/4)^2 / Pi*(1/2)^2 = 1/4

So the probability the first shot is NOT in the small circle is 1 - 1/4 = 3/4. Thus the probability all five shots are NOT in the small circle is (3/4)^5, and subtracting this from 1, our answer is 1 - (3/4)^5.
For online GMAT math tutoring, or to buy my higher-level Quant books and problem sets, contact me at ianstewartgmat at gmail.com

ianstewartgmat.com
Join the discussion